# Two-sample comparison through additive tree models for density ratios

### v1

@@ C.3 Notes on the implementation of the density ratio estimators
In the proposed boosting algorithms, we prune leaf nodes that include only observations from one group to avoid division by zero in computing the optimal values of leaf node parameters.
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross validation based on the classification error reported by the function. We found that the selected number tends to be close to zero when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10.
@@ Appendix D Additional figures
We provide additional visualizations of the results for the settings described in Section 4.1. Figure 8 shows the estimated log-density ratios for the first (global shift) and third (local dispersion) scenarios under the balanced sample size setting. The estimated ratios for the second (local shift) scenario under the unbalanced sample size setting are displayed in Figure 9.
Figure 8: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.
Figure 9: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.

### v2

@@ C.3 Notes on the implementation of the density ratio estimators
In the proposed boosting algorithms, we prune leaf nodes that contain observations from only one group to avoid division by zero in computing the optimal values of leaf node parameters.
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross-validation based on the classification error reported by the function. We found that the selected number tends to be close to zero when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10 manually.
@@ Appendix D Additional figures
We provide additional visualizations corresponding to the settings described in Section 4.1. Figure 8 shows the estimated log-density ratios for the first (global shift) and third (local dispersion difference) scenarios under the balanced sample size setting, and Figure 9 shows the estimated ratios for the second (local shift) scenario under the unbalanced sample size configuration.
Figure 8: The estimated log-density ratios obtained using the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
Figure 9: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.